# Emergent symmetries in prethermal phases of periodically driven quantum systems

### v1

@@ Abstract
Periodically driven closed quantum systems are expected to eventually heat up to infinite temperature reaching a steady state described by a circular orthogonal ensemble (COE). However, such finite driven systems may exhibit sufficiently long prethermal regimes; their properties in these regimes are qualitatively different from that in their infinite temperature steady states. These, often experimentally relevant, prethermal regimes host a wide range of phenomena; they may exhibit dynamical localization and freezing, host Floquet scars, display signatures of Hilbert space fragmentation, and exhibit time crystalline phases. Such phenomena are often accompanied by emergent approximate dynamical symmetries which have no analogue in equilibrium systems. In this review, we provide a pedagogical introduction to the origin and nature of these symmetries and discuss their role in shaping the prethermal phases of a class of periodically driven closed quantum systems.
@@ I Introduction
The physics of closed quantum systems driven out of equilibrium has been a subject of intense research in recent years both theoretically [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] and experimentally using several ultracold atom [12, 13, 14, 15, 16, 17, 18], ion trap [19] and superconducting qubit [20] platforms. Initial endeavors in this direction constituted study of quench and ramp protocols [1, 2]. In the former case, a parameter $\lambda$ of the Hamiltonian $H[\lambda]$ governing the dynamics system is abruptly changed at $t=0$ from $\lambda_{i}$ to $\lambda_{f}$ and the subsequent evolution of the initial state (often taken as an eigenstate $H[\lambda_{i}]$) is studied [21, 23, 24, 25, 26]. In the latter class of drive protocols, $\lambda$ is varied continually as a function of time with a fixed rate $\tau^{-1}$: $\lambda(t)=\lambda_{i}f(t/\tau)$, where $f$ is chosen such that $f(t_{i}/\tau)=1$ and $f(t_{f}\tau)=\lambda_{f}/\lambda_{i}$ [3, 27, 28, 29, 30, 31]. In both cases, several quantities of the system such as excitation (defect) density $n_{d}$ , excess (residual) energy $Q$, correlation functions $C$, and entanglement entropies $S$ (Renyi and von-Neumann) are studied as a function of time; these quantities provided information about spread of the state of the system.
@@
The study of such driven systems in cases where the Hamiltonian moves across or passes through a quantum critical point at $\lambda=\lambda_{c}$ has been of interest; for slow ramp protocols, $n_{d}$ and $Q$ are known to scale with ramp rate $\tau^{-1}$ with universal exponents characterized by the dynamical critical exponent $z$ and the correlation length exponent $\nu$ of the critical point. This phenomenon is known as Kibble-Zurek scaling [1, 2, 3, 4, 32, 33, 34, 35, 36, 37, 28, 38, 39, 40, 41]. A similar scaling phenomenon, obtained through study of fidelity $F(\delta\lambda;t)=\ln|\langle\psi(\lambda_{i})|\psi(\lambda+\delta\lambda;t)\rangle|$ for $\delta\lambda=\lambda-\lambda_{i}$; the scaling is observed fidelity susceptibility $\chi=1-F$: $\chi\sim|\delta\lambda|^{\alpha}$, where $\alpha$ is once again determined by the universality of the critical point [25].
@@
The later studies of non-equilibrium dynamics concentrated on periodic drive protocols [3, 4, 5, 6, 7, 8, 9, 10, 11]. This shift of focus occurred due to several reasons. First, the evolution operator $U$ of a periodically driven system characterized by a drive period $T$, can be described at stroboscopic times $t=nT$ (where $n$ is an integer) by a time-independent Floquet Hamiltonian $H_{F}$: $U(nT,0)=\exp[-iH_{F}nT]$ [8, 9, 10, 42]. This provides a large simplification; several properties of the driven system can be inferred from that of a time independent Hamiltonian $H_{F}$. Although it is not usually feasible to have an exact analytical expression of $H_{F}$, several perturbative schemes exists which can provide analytical, albeit qualitative, understanding of its properties. These perturbative schemes are specially useful in the high drive frequency and/or drive amplitude regimes; examples of such schemes include Magnus expansion and Floquet perturbation theory (Magnus expansion in a rotated frame) [11, 43, 44]. In contrast, for low frequencies, such techniques are unavailable for a generic non-integrable driven quantum system. The convergence of the perturbative schemes such as Magnus expansion in lower and intermediate drive frequency range is also an yet unresolved issue [45, 46, 47]. In the high drive frequency or amplitude regime where Magnus expansion or Floquet perturbation theory converges, the Floquet Hamiltonian can typically be approximated by a local Hamiltonian. In contrast, in the low-frequency regime, it becomes non-local. Nevertheless, even in the this regime, numerical techniques such as exact diagonalization (ED) provide us with knowledge of the Floquet spectrum specially for finite systems in $d=1$; this allows one to analyze the property of such driven systems using a time-independent Hamiltonian.
@@
Second, it has been shown that such driven systems often exhibits a prethermal regime before eventually heating up to infinite temperature [48, 49]. This prethermal timescale can be very large in the large drive amplitude/frequency regime; consequently, they become experimentally relevant. For example in experiments using ultracold atom platforms [12, 13, 14, 15, 16, 17, 18], which are reasonably well isolated from the environment, have a finite system lifetime [12]; it turns out that in certain cases the prethermal regime of the driven system can match or exceed this lifetime making it the singularly important regime which can be probed in these experiments. In such a prethermal regime, the driven system is well approximated by a local Floquet Hamiltonian and exhibits several phenomena which have no analogue in undriven quantum systems. These include dynamical freezing [50, 51, 52, 53, 54, 55, 56, 57, 59, 58, 60, 61, 62, 89], dynamical localization [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], signatures of prethermal Hilbert space fragmentation (HSF) [79, 80, 81, 82], presence of Floquet scars [83, 84, 85, 86, 87, 88, 89], and prethermal time crystalline phases[90, 91, 92, 93, 94, 95, 96, 97]; a detailed analysis of such phenomena has been a subject of major endeavor in recent years.
@@
In this review, we shall provide a pedagogical introduction to the properties of some such prethermal phases. Often, as we shall discuss, these phases are accompanied by emergent approximate symmetries. These symmetries play a crucial role in shaping the nature of these prethermal phases. The aim of this review will be to highlight the role of these emergent symmetries and discuss their origin. In most cases which will be discussed in this review, the properties of the prethermal phases will be shown to be controlled by a local, perturbative Floquet Hamiltonian, $H_{F}^{(1)}$, which can be analytically obtained as the lowest order term in some properly chosen perturbation expansion. It turns out that at special drive frequencies or amplitudes, $H_{F}^{(1)}$ may respect additional symmetries or impose additional dynamical constraints; these symmetries or constraints are approximate in the sense that they are not respected by the exact Floquet Hamiltonian $H_{F}$. However, in the large drive amplitude or frequency regime, where $H_{F}^{(1)}$ controls the evolution of the system up to a large prethermal time, such emergent symmetries or constraints shape the dynamics of the systems. In what follows, we shall explain this phenomenon in several contexts. The role of such emergent symmetries behind dynamical freezing in both Hermitian and non-Hermitian driven systems shall be discussed in Sec. II. This will be followed, in Sec. III by a discussion of Floquet scars and prethermal HSF. Next in Sec. IV, we discuss the phenomenon of prethermal time crystals . Finally, we discuss some open issues and conclude in Sec. V.
@@ II Dynamical freezing and localization
In this section, we shall introduce the concept of emergent approximate symmetries using one of the simplest possible setup, namely, a periodically driven Ising model in a transverse field in $d=1$. We shall discuss, in Sec. II.1, how such a driven system exhibits prethermal dynamical freezing. In Sec. II.2, we shall analyze another setup, namely, interacting 1D fermions in the presence of electric field, which exhibits dynamical localization due to a similar approximate emergent symmetry.
@@ II.1 Dynamical Freezing
The $d=1$ Ising model in a transverse field is a prototypical model for studying quantum dynamics in both Hermitian and non-Hermitian contexts [98, 99]. The Hamiltonian of this model, in terms of interacting spins on a 1D chain, is given by
$H_{\rm Ising}$ $=$ $-J\sum_{\langle\ell j\rangle}\sigma_{\ell}^{x}\sigma_{j}^{x}-h_{0}\sum_{j}\sigma_{j}^{z}$ (1)
where $J$ is the interaction amplitude, $h_{0}$ is the strength of the transverse field in units of energy, $\sigma_{j}^{\alpha=x,z}$ denotes usual Pauli matrices on site $j$ of the chain, and $\langle\ell j\rangle$ denotes sum over sites such that the site $j$ is one of the nearest neighbors of the site $\ell$. The transverse field $h_{0}$ is a real number for the Hermitian model; for the non-Hermitian Ising model $h_{0}=h+i\gamma$ is a complex number and $\gamma$ denotes the strength of dissipation [99].
We shall first consider the driven non-Hermitian model for which $h_{0}$ is a complex parameter; the results for the Hermitian case following by taking the limit ${\rm Im}[h_{0}]=0$. The drive is implemented by making $h_{0}$ time dependent; in what follows, we shall use
@@
$\tau_{j}^{+(-)}$ $=$ $\left(\prod_{\ell=1}^{j-1}-\tau_{\ell}^{z}\right)f_{j}^{\dagger}(f_{j}),\quad\tau_{j}^{z}=2f_{j}^{\dagger}f_{j}-1$ (3)
@@
$f_{j}=\frac{1}{\sqrt{N}}\sum_{k\in\text{BZ}}e^{i\pi/4}e^{-ikj}\hat{f}_{k}$ (4)
@@
$H_{k}$ $=$ $\tau_{z}(h_{0}(t)-\cos ka+i\gamma)+\tau_{x}\sin ka$
where $\vec{\tau}=(\tau_{x},\tau_{y},\tau_{z})$ are standard Pauli matrices in particle-hole space of the fermions, $J$ is set to unity, $a$ is the lattice spacing, ${\rm BZ}/2$ denotes positive half of the Brillouin zone, and $\tau^{\pm}=\tau_{x}\pm i\tau_{y}$.
@@
$U_{0}(t,0)$ $=$ $\prod_{k}U_{0k}(t,0)$
$U_{0k}(t,0)$ $=$ $e^{-2ih_{1}\sin\omega_{d}t\tau_{z}/(\hbar\omega_{D})}.$ (6)
@@
We note that $U_{0}$ has a single Pauli matrix which allows one to take care of the time ordering easily. Furthermore, $U_{0}(T,0)=I$ (where $I$ denotes the identity matrix) leading to $H_{F}^{(0)}=0$.
@@
Next one obtains the higher order terms in $U(T,0)$. The first order perturbative correction to the evolution operators is obtained as [60, 61, 11, 43, 44]
$U_{1}(T,0)$ $=$ $\prod_{k}\left(\frac{-i}{\hbar}\right)\int_{0}^{T}U_{0k}^{\dagger}(t,0)H^{\prime}_{k}U_{0k}(t,0)$ (7)
$H^{\prime}_{k}$ $=$ $2(\tau_{z}(h_{s}-\cos ka+i\gamma)+\tau_{x}\sin ka)$
@@
This expression can be evaluated in a straightforward manner and one obtains $H_{F}^{(1)}=i\hbar U_{1}(T,0)/T$ given by [60, 61]
$H_{F}^{(1)}$ $=$ $2\sum_{k\in{BZ/2}}\Big{[}\tau_{z}(h_{s}-\cos ka+i\gamma)$ (8)
$+\tau_{x}J_{0}(\mu)\sin ka\Big{]}$
@@
where $J_{n}$ denotes the $n^{\rm th}$ order Bessel function of the first kind and $\mu=4h_{1}/(\hbar\omega_{d})$ The first term in Eq. 8 follows from the fact that the first term of $H^{\prime}_{k}$ commutes with $U_{0k}$. For obtaining the second term, one needs to note that action of $\tau_{x}$ on an eigenstate of $\tau_{z}$ simply flips a pseudospin; the details of the calculation is presented in Refs. 60, 61. The second order term in the perturbative expansion can also be obtained in a similar manner. A somewhat cumbersome calculation, outlined in Refs. 60, 61, leads to
$H_{F}^{(2)}=\sum_{k}\Big{[}-4\tau_{z}\sin^{2}ka\sum_{n=0}^{\infty}\frac{J_{0}(\mu)J_{2n+1}(\mu)}{(2n+1)\hbar\omega_{D}}$
$+\tau_{x}\sin ka(h_{s}-\cos ka+i\gamma)\sum_{n=0}^{\infty}\frac{4J_{2n+1}(\mu)}{(2n+1)\hbar\omega_{D}}\Big{]}$ (9)
@@
We note that $H_{F}^{(2)}$ is suppressed by a factor of $1/\omega_{D}$ compared to $H_{F}^{(1)}$ indicating the efficacy of the perturbative expansion at high drive frequencies.
@@
In what follows we are going to analyze the properties of the system using the expressions of $H_{F}^{(1)}$ and $H_{F}^{(2)}$. To this end, we first set $\gamma=0$ and focus on the Hermitian Ising chain. From $H_{F}^{(1)}$ we note that at special frequencies $\omega_{D}=\omega_{n}^{\ast}$, for which $J_{0}(\mu)=0$; this leads to
$\frac{4h_{1}}{\hbar\omega_{n}^{\ast}}=\beta_{n},$ (10)
@@
where $\beta_{n}$ denotes $n^{\rm th}$ zero of $J_{0}(\mu)$. At these drive frequencies, the second term in Eq. 8 vanishes; consequently $H_{F}^{(1)}$ commutes with $\tau_{z}$. This leads to an additional approximate emergent symmetry resulting in conservation of the fermion density $n_{k}=\psi_{k}^{\dagger}\tau_{z}\psi_{k}$ (which is equivalent to conservation of the magnetization $M_{z}=\sum_{j}\sigma_{j}^{z}$ in the original spin model) at stroboscopic times $t=nT$ where $n$ is an integer. This emergent symmetry is approximate since it is broken by $H_{F}^{(2)}$; this is easily seen from the second term of Eq. 9 which does not commute with $\tau_{z}$ at these frequencies. We note here that such emergent symmetries occur for other drive protocols; however, the condition given by Eq. 10 for their occurrence changes. For example, for a square-pulse drive protocol which has been studied extensively in the literature in this context [58, 60, 61, 79], yields the condition to be $2h_{1}/(\hbar\omega_{n}^{\ast})=n$ for integer $n$.
@@
Figure 1: (a) Plot of $M_{z}$ as a function of number of drive cycles $n$ showing dynamical freezing at special drive frequency $\omega_{D}=\omega_{2}^{\ast}=4h_{1}/\beta_{2}$ and evolution to its steady state value away from the special frequency (red curve). The inset shows $M_{z}$ averaged over $n_{0}=200$ cycles (from $n=1800$ to $n=2000$) as a function of $h_{1}$ at $\omega_{D}=\omega_{1}^{\ast}$ indicating loss of freezing at low drive amplitude. For all plots $h_{1}=20$, $h_{s}=0.1$, $L=500$, and $J=1$. (b) Plot of $\alpha$ (where $S_{L/2}\sim\alpha\ln L$) as a function of $\omega_{D}$ and $\gamma$ for a cosine drive protocol. The plot, adapted from Ref. 61, clearly indicates that $\alpha$ vanishes at special drive frequencies $\omega_{n}^{\ast}$ indicated in the figure. For all plots $h_{1}=20J$, $h_{s}=0.1$, $J=1$ and $L=1000$. See text for details.
@@
The consequence of such an emergent symmetry leading to dynamical freezing of the magnetization $M_{z}$ (or equivalently $\sum_{k}\langle n_{k}\rangle$) at stroboscopic times $t=nT$ to its initial value up to a prethermal timescale where $H_{F}^{(1)}$ dominates the dynamics (see Fig. LABEL:sec2en(a)). This phenomenon is first noted in Ref. 50; there are several subsequent works on this phenomenon which confirmed this behavior [51, 52, 53, 58] (see Fig. 1(a)). This freezing is approximate; the dynamics of $M_{z}$ is eventually restored when effects of $H_{F}^{(2)}$ and other higher order terms becomes important.

### v2

@@ Abstract
Periodically driven closed quantum systems are expected to eventually heat up to infinite temperature ; reaching a steady state described by a circular orthogonal ensemble (COE). However, such finite driven systems may exhibit sufficiently long prethermal regimes; their properties in these regimes are qualitatively different from that of their corresponding infinite temperature steady states. These, often experimentally relevant, prethermal regimes host a wide range of phenomena; they may exhibit dynamical localization and freezing, host Floquet scars, display signatures of Hilbert space fragmentation, and exhibit time crystalline phases. Such phenomena are often accompanied by emergent approximate dynamical symmetries which have no analogue in equilibrium systems. In this review, we provide a pedagogical introduction to the origin and nature of these symmetries and discuss their role in shaping the prethermal phases of a class of periodically driven closed quantum systems.
@@ I Introduction
The physics of closed quantum systems driven out of equilibrium has been a subject of intense research in recent years both theoretically rev1 ; rev2 ; rev3 ; rev4 ; rev5 ; rev6 ; rev7 ; rev8 ; rev9 ; rev10 ; rev11 and experimentally using several ultracold atom rev12 ; exp0 ; exp1 ; exp2 ; exp3 ; exp4 ; exp5 , ion trap revion and superconducting qubit revqu platforms. Initial endeavors in this direction constituted studies involving quench and ramp protocols rev1 ; rev2 . In the former case, a parameter $\lambda$ of the Hamiltonian $H[\lambda]$ governing the dynamics system is abruptly changed at $t=0$ from $\lambda_{i}$ to $\lambda_{f}$; the subsequent evolution of the initial state (often taken as an eigenstate of $H[\lambda_{i}]$) is studied calabrese1 ; subir1 ; dsen0 ; anatoly0 ; sdas0 . In the latter class of drive protocols, $\lambda$ is varied continually as a function of time with a fixed rate $\tau^{-1}$: $\lambda(t)=\lambda_{i}f(t/\tau)$. Usually the function $f$ is chosen such that $f(t_{i}/\tau)=1$ and $f(t_{f}/\tau)=\lambda_{f}/\lambda_{i}$ rev3 ; sdas1 ; sdas2 ; rajdeep1 ; rajdeep2 ; trip1 . In both cases, several quantities of the system such as excitation (defect) density $n_{d}$ , excess (residual) energy $Q$, correlation functions $C$, and entanglement entropies $S$ (Renyi and von-Neumann) are studied as a function of time.
@@
The study of such driven systems in cases where the Hamiltonian moves across or passes through a quantum critical point at $\lambda=\lambda_{c}$ has been of particular interest. For slow ramp protocols, $n_{d}$ and $Q$ are known to scale with ramp rate $\tau^{-1}$ with universal exponents characterized by the dynamical critical exponent $z$ and the correlation length exponent $\nu$ of the critical point. This phenomenon is known as Kibble-Zurek scaling rev1 ; rev2 ; rev3 ; rev4 ; kibble1 ; zurek1 ; anatoly1 ; anatoly2 ; dsen1 ; dsen2 ; sdas2 ; sdas3 ; sdas4 ; adutta1 ; rajdeep3 . A similar scaling phenomenon was obtained through study of fidelity $F(\delta\lambda;t)=\ln|\langle\psi(\lambda_{i})|\psi(\lambda_{i}+\delta\lambda;t)\rangle|$ for $\delta\lambda=\lambda-\lambda_{i}$; the scaling of the fidelity susceptibility $\chi=1-F$ has been found to be $\chi\sim|\delta\lambda|^{\alpha}$, where $\alpha$ is once again determined by the universality of the critical point anatoly0 .
@@
The later studies of non-equilibrium dynamics concentrated on periodic drive protocols rev3 ; rev4 ; rev5 ; rev6 ; rev7 ; rev8 ; rev9 ; rev10 ; rev11 . This shift of focus occurred due to several reasons. First, the evolution operator $U$ of a periodically driven system, characterized by a drive period $T$, can be described at stroboscopic times $t=nT$ (where $n$ is an integer) by a time-independent Floquet Hamiltonian $H_{F}$: $U(nT,0)=\exp[-iH_{F}nT]$ rev8 ; rev9 ; rev10 ; flref1 . This provides a large simplification; several properties of the driven system can be inferred from that of a time independent Hamiltonian $H_{F}$. Although it is not usually feasible to have an exact analytical expression of $H_{F}$, several perturbative schemes exists which can provide analytical, albeit qualitative, understanding of its properties. These perturbative schemes are specially useful in the high drive frequency and/or drive amplitude regimes; examples of such schemes include Magnus expansion and Floquet perturbation theory (Magnus expansion in a rotated frame) rev11 ; aa1 ; nc1 . In contrast, for low frequencies, such techniques are unavailable for a generic non-integrable driven quantum system. The convergence of the perturbative schemes such as Magnus expansion in lower and intermediate drive frequency range is also an yet unresolved issue mori1 ; pre2 ; anatoly3 ; abanin1 . In the high drive frequency or amplitude regime where Magnus expansion or Floquet perturbation theory converges, the Floquet Hamiltonian can typically be approximated by a local Hamiltonian. In contrast, in the low-frequency regime, it becomes non-local. Nevertheless, even in the this regime, numerical techniques such as exact diagonalization (ED) provide us with knowledge of the Floquet spectrum specially for finite systems in $d=1$; this allows one to analyze the properties of such driven systems using a time-independent Hamiltonian.
@@
Second, it has been shown that such driven systems often exhibit a prethermal regime before eventually heating up to infinite temperature prethermrev1 ; prethermrev2 . This prethermal timescale can be very large in the large drive amplitude/frequency regime; consequently, they become experimentally relevant. For example, in experiments using ultracold atom platforms rev12 ; exp0 ; exp1 ; exp2 ; exp3 ; exp4 ; exp5 , which are reasonably well isolated from the environment, have a finite system lifetime rev12 ; it turns out that in certain cases the prethermal regime of the driven system can match or exceed this lifetime. This allows one to probe the prethermal regime in these experiments. In such a prethermal regime, the driven system is well approximated by a local Floquet Hamiltonian. Moreover, it exhibits several phenomena which have no analogue in undriven quantum systems. These include dynamical freezing adas1 ; adas2 ; adas3 ; pekker1 ; deb1 ; uma1 ; camilo1 ; apal1 ; koch1 ; adasnew ; tb1 ; tb2 ; huo1 ; liu1 , dynamical localization dynloc1 ; dynloc2 ; dynloc3 ; dynloc4 ; dynloc5 ; tanay1 ; fava1 ; rg1 ; galit1 ; martinez1 ; guoexp , signatures of prethermal Hilbert-space fragmentation sg1 ; sg2 ; xu1 ; zhang1 , presence of Floquet scars pretko1 ; bm1 ; mituza1 ; bm2 ; lukinsc1 ; papic1 ; liu1 , and prethermal time crystalline phasestcrev1 ; tcrev2 ; tcrev3 ; tcrev4 ; tcrev5 ; tcpap1 ; tcpap2 ; tcpap3 . A detailed analysis of such phenomena has been a subject of intense research in recent years.
@@
In this review, we shall provide a pedagogical introduction to the properties of some such prethermal phases. Often, as we shall discuss, these phases are accompanied by emergent approximate symmetries. These symmetries play a crucial role in shaping the nature of these prethermal phases. The aim of this review will be to highlight the role of these emergent symmetries and discuss their origin. In most cases which will be discussed in this review, the properties of the prethermal phases will be shown to be controlled by a local, perturbative Floquet Hamiltonian, $H_{F}^{(1)}$, which can be analytically obtained as the lowest order term in some properly chosen perturbation expansion.
It turns out that for several periodically driven systems, $H_{F}^{(1)}$ may respect additional symmetries or impose additional dynamical constraints at special drive frequencies or amplitudes. These symmetries or constraints are approximate in the sense that they are not respected by the exact Floquet Hamiltonian $H_{F}$. However, in the large drive amplitude or frequency regime, where $H_{F}^{(1)}$ controls the evolution of the system up to a large prethermal time, such emergent symmetries or constraints shape the dynamics of these driven systems.
In what follows, we shall explain this phenomenon in several contexts. The role of such emergent symmetries behind dynamical freezing in both Hermitian and non-Hermitian driven systems shall be discussed in Sec. II. This will be followed, in Sec. III by a discussion of Floquet scars and prethermal Hilbert-space fragmentation. Next in Sec. IV, we discuss the phenomenon of prethermal time crystals . Finally, we discuss some open issues and conclude in Sec. V.
@@ II Dynamical freezing and localization
In this section, we shall introduce the concept of emergent approximate symmetries using one of the simplest possible setup, namely, a periodically driven Ising model in a transverse field in $d=1$. We shall discuss, in Sec. II.1, how such a driven system exhibits prethermal dynamical freezing. In Sec. II.2, we shall analyze another setup, namely, interacting 1D fermions in the presence of electric field, which exhibits dynamical localization due to a similar approximate emergent symmetry. We note that both these phenomena involve drive-induced localization. While the former is typically used to describe localization in the Fock space, the latter refers to localization in the real space.
@@ II.1 Dynamical Freezing
The $d=1$ Ising model in a transverse field is a prototypical model for studying quantum dynamics in both Hermitian and non-Hermitian contexts subirbook ; lec1 ; nhrev . The Hamiltonian of this model, in terms of interacting spins on a 1D chain, is given by
$H_{\rm Ising}$ $=$ $-J\sum_{\langle\ell j\rangle}\sigma_{\ell}^{x}\sigma_{j}^{x}-h_{0}\sum_{j}\sigma_{j}^{z},$ (1)
where $J$ is the interaction amplitude, $h_{0}$ is the strength of the transverse field in units of energy, $\sigma_{j}^{\alpha=x,z}$ denotes usual Pauli matrices on site $j$ of the chain, and $\langle\ell j\rangle$ denotes sum over sites such that the site $j$ is one of the nearest neighbors of the site $\ell$. The transverse field $h_{0}$ is a real number for the Hermitian model; for the non-Hermitian Ising model $h_{0}=h+i\gamma$ is a complex number and $\gamma$ denotes the strength of dissipation nhrev . We note here that the non-Hermitian transeverse field Ising model is a prototypical non-Hermitian model. It constitutes a representation of an open Ising chain in the presence of measurement in the so-called no-click limit nhrev . In this limit, the induced non-Hermiticity in the system’s Hamiltonian can be thought of as an effect of the measurement back-action on the system’s degrees of freedom. Its analysis, as we discuss below, allows us to relate the steady state properties of the driven chain to emergent approximate symmetries; this feature has no counterpart in Hermitian quantum models.
We shall first consider the driven non-Hermitian model for which $h_{0}$ is a complex parameter. This will be followed by the results for the Hermitian case taking the limit ${\rm Im}[h_{0}]=0$. The drive is implemented by making $h_{0}$ time dependent; in what follows, we shall use
@@
$\tau_{j}^{+(-)}$ $=$ $\left(\prod_{\ell=1}^{j-1}-\tau_{\ell}^{z}\right)f_{j}^{\dagger}(f_{j}),\quad\tau_{j}^{z}=2f_{j}^{\dagger}f_{j}-1,$ (3)
@@
$f_{j}=\frac{1}{\sqrt{N}}\sum_{k\in\text{BZ}}e^{i\pi/4}e^{-ikj}\hat{f}_{k},$ (4)
@@
$H_{k}$ $=$ $\tau_{z}(h_{0}(t)-\cos ka+i\gamma)+\tau_{x}\sin ka.$
Here $\vec{\tau}=(\tau_{x},\tau_{y},\tau_{z})$ are standard Pauli matrices in particle-hole space of the fermions, $J$ is set to unity, $a$ is the lattice spacing, ${\rm BZ}/2$ denotes the positive half of the Brillouin zone, and $\tau^{\pm}=\tau_{x}\pm i\tau_{y}$.
@@
$U_{0k}(t,0)$ $=$ $e^{-2ih_{1}\sin\omega_{d}t\tau_{z}/(\hbar\omega_{D})},$ (6)
@@
for a momentum mode $k$. We note that $U_{0k}$ has a single Pauli matrix which allows one to take care of the time ordering easily. Furthermore, $U_{0k}(T,0)=I$ (where $I$ denotes the identity matrix) leading to $H_{Fk}^{(0)}=0$ for all $k$.
@@
Next one obtains the higher order terms in $U(T,0)$. The first order perturbative correction to the evolution operators is obtained as tb1 ; tb2 ; rev11 ; aa1 ; nc1
[…]
$H^{\prime}_{k}$ $=$ $2(\tau_{z}(h_{s}-\cos ka+i\gamma)+\tau_{x}\sin ka)$ (7)
@@
This expression can be evaluated in a straightforward manner and one obtains $H_{Fk}^{(1)}=i\hbar U_{1k}(T,0)/T$ given by tb1 ; tb2
$H_{Fk}^{(1)}$ $=$ $2\Big{[}\tau_{z}(h_{s}-\cos ka+i\gamma)$
$+\tau_{x}J_{0}(\mu)\sin ka\Big{]},$
$H_{F}^{(1)}$ $=$ $\sum_{k}\psi_{k}^{\dagger}H_{Fk}^{(1)}\psi_{k}.$ (8)
@@
Here $J_{p}$ denotes the $p^{\rm th}$ order Bessel function of the first kind and $\mu=4h_{1}/(\hbar\omega_{d})$. The first term in Eq. 8 follows from the fact that the first term of $H^{\prime}_{k}$ commutes with $U_{0k}$. For obtaining the second term, one needs to note that the action of $\tau_{x}$ on an eigenstate of $\tau_{z}$ simply flips a pseudospin. The details of the relevant calculation is starightforward and presented in Refs. tb1, ; tb2, . The second order term in the perturbative expansion can also be obtained in a similar manner. A somewhat cumbersome calculation, outlined in Refs. tb1, ; tb2, , leads to
$H_{Fk}^{(2)}=\Big{[}-4\tau_{z}\sin^{2}ka\sum_{n=0}^{\infty}\frac{J_{0}(\mu)J_{2n+1}(\mu)}{(2n+1)\hbar\omega_{D}}$
$+\tau_{x}\sin ka(h_{s}-\cos ka+i\gamma)\sum_{n=0}^{\infty}\frac{4J_{2n+1}(\mu)}{(2n+1)\hbar\omega_{D}}\Big{]}.$
$H_{F}^{(2)}=\sum_{k\in\text{BZ}/2}\psi_{k}^{\dagger}H_{Fk}^{(2)}\psi_{k}.$ (9)
@@
We note that $H_{Fk}^{(2)}$ is suppressed by a factor of $1/\omega_{D}$ compared to $H_{Fk}^{(1)}$ indicating the efficacy of the perturbative expansion at high drive frequencies.
@@
In what follows we are going to analyze the properties of the system using the expressions of $H_{Fk}^{(1)}$ and $H_{Fk}^{(2)}$. To this end, we first set $\gamma=0$ and focus on the Hermitian Ising chain. From $H_{Fk}^{(1)}$ we note that at special frequencies $\omega_{D}=\omega_{m}^{\ast}$, for which $J_{0}(\mu)=0$; this leads to
$\frac{4h_{1}}{\hbar\omega_{m}^{\ast}}=\beta_{m},$ (10)
@@
where $\beta_{m}$ denotes $m^{\rm th}$ zero of $J_{0}(\mu)$. At these drive frequencies, the second term in Eq. 8 vanishes; consequently $H_{Fk}^{(1)}$ commutes with $\tau_{z}$. This leads to an additional approximate emergent symmetry resulting in conservation of the fermion density $n_{k}=\psi_{k}^{\dagger}\tau_{z}\psi_{k}$ (which is equivalent to conservation of the magnetization $M_{z}=\sum_{j}\sigma_{j}^{z}$ in the original spin model) at stroboscopic times $t=nT$ where $n$ is an integer. This emergent symmetry is approximate since it is broken by $H_{Fk}^{(2)}$; this is easily seen from the second term of Eq. 9 which does not commute with $\tau_{z}$ at these frequencies. We note here that such emergent symmetries occur for other drive protocols; however, the condition given by Eq. 10 for their occurrence changes. For example, a square-pulse drive protocol, which has been studied extensively in the literature in this context adasnew ; tb1 ; tb2 ; sg1 , leads to a condition $2h_{1}/(\hbar\omega_{m}^{\ast})=m$ for integer $m$.
@@
Figure 1: Left Panel: Plot of $M_{z}$ as a function of number of drive cycles $n$ showing dynamical freezing at special drive frequency $\omega_{D}=\omega_{2}^{\ast}=4h_{1}/\beta_{2}$ and evolution to its steady state value away from the special frequency (red curve). The inset shows $M_{z}$ averaged over $n_{0}=200$ cycles (from $n=1800$ to $n=2000$) as a function of $h_{1}$ at $\omega_{D}=\omega_{1}^{\ast}$ indicating loss of freezing at low drive amplitude. For all plots $h_{1}=20$, $h_{s}=0.1$, $L=500$, and $J=1$. Right Panel: Plot of $\alpha$ (where $S_{L/2}\sim\alpha\ln L$) as a function of $\omega_{D}$ and $\gamma$ for a cosine drive protocol. The plot, adapted from Ref. tb2, , clearly indicates that $\alpha$ vanishes at special drive frequencies $\omega_{n}^{\ast}$ indicated in the figure. For all plots $h_{1}=20J$, $h_{s}=0.1$, $J=1$ and $L=1000$. See text for details.
@@
The consequence of such an emergent symmetry is dynamical freezing of the magnetization $M_{z}$ (or equivalently $\sum_{k}\langle n_{k}\rangle$) at stroboscopic times $t=nT$ up to a prethermal timescale where $H_{F}^{(1)}$ dominates the dynamics (left panel of Fig. 1). This phenomenon was first noted in Ref. adas1, ; there are several subsequent works on this phenomenon which confirmed this behavior adas2 ; adas3 ; pekker1 ; adasnew (left panel of Fig. 1). Such a freezing is approximate; the dynamics of $M_{z}$ is eventually restored when effects of $H_{F}^{(2)}$ and other higher order terms becomes important.